Fermi National Laboratory
 

Accelerator Update for August 6, 2007 - August 10, 2007

 

Fermilab Shutdown Report

LINAC

  • Low energy Low Level RF upgrade – ~ 5% complete
  • H- Vacuum upgrade – 50% complete

Booster

Electrical Work:

  • 480/120V connection to the racks at BWT complete
  • Cable pull progressing well (~ 45% complete)

Water work:

LCW water work in the Booster West Gallery

  • Booster West Gallery isolated from CUB
  • All RF stations have been removed out the way
  • Prep work and set up for contractors almost complete
  • Tunnel water work had a great start and is going well
  • Replacing black tubing with Kynar (~60% complete)

BE AWARE:

      • 4” and 6” copper pipe are staged in the gallery.

Corrector Installation:

  • Disassemble prototype from L4 and shipped back to MTF for tests
  • 4 fully assembled magnets are ready for installation
  • One of them has been installed at L19
  • Today’s plan is to install another one at L17

Next week:

  • Finish the cable pull
  • Contractors will start installation of the 4” and 6” pipe. There will be a lot of activity with parts in the hallway.
  • Start installing the filtration system
  • Align both correctors and continue prepping other location in preparation for installation.

Antiproton Source

Mechanical

  • Leak Check (4)
    • Finished checking D/A & A40
      • D:ESEP fixed
  • D:ISEP Shielding
    • Cutting steel
  • Ion Chamber removal
  • Deb DCCT motorized stand
    • Will move GAPMON?
  • D:EKIK ceramic tube
  • AP2 Water Hose Replacement
  • Leaking LQA
    • Disconnected; as-found survey
  • Replace leaking Debuncher Sextuple

Cryo

  • Valve Stem replaced
  • Leak Check (2)
  • LN2 transfer line maintenance

Target Station

  • Water Systems (2)
  • Temperature Monitoring (2)
  • Target Rotation
  • Storage
  • SEM Module
  • LVDTs
  • Sample Retrieval
  • Lens motion
  • Air Exhaust Stack
  • 10mm-2 into transformer
    • Prepping
  • 10mm-3 into transformer
    • Prepping
  • Lens replacement

Cooling Systems

  • Deb 3&4 Trans Notch Filters
    • Work in progress
  • Deb Double Notch Filters
    • Work in progress
  • Stacktail 4-8GHz Equalizer
    • Work in progress
  • TWT balance & trip points
  • Stacktail Filter 3 cable change

Electrical (Pbar & EE Support)

  • Shunt Move
  • Shunt Repairs (6)
  • PS Maintenance
  • Lens SCRs
    • Tested; installing
  • Regulation Electronics
  • Ion Pump maintenance
  • Deb Shunt re-arrangement
    • Await plan

Other

  • Interlocks/Safety System (3)
  • Repair cathodic protection
  • Dirt holes under buildings
  • DCCT calibration
  • Baking System checkout
  • RF & TWT filters maintenance
  • Lights Checkout
    • FESS to fix
  • Acc BPM repairs
  • Activation Survey
    • Finished AP1 & Accumulator
  • Inductance checkout
  • PM ODH Fans & Sumps
  • Tunnel Wireless
  • Install Leak Detector
  • SEM maintenance

Main Injector

Began installation of M30 collimator


Miscellaneous maintenance

Recycler

Pelletron

  • Open since Tuesday (8/7)
    • Ron did not find anything abnormal
    • Usual amount of chain dust; no shaft dust
    • Found only two needles (less efficient in high voltage regulation)
    • Needles will be replaced
    • Measured GVM face to terminal shell distance.
    • Removal black poly dust
    • Chain: clean or replace?
    • New DCCT section installed, leak checked and under vacuum since 9 AM Wednesday.
      • Ready for alignment
    • V303 move planned; move qt303 DS and move V303 to where qt303 was.
    • Trombone additional quadrupoles installation issues
      • Under way starting at 601
      • Need to move MI buss support at 604
      • MI RF PA’s nearby
      • Need to talk to HLRF about stations 14 and 18

Tevatron

Things To Come

  • E1 and E0 warm up and leak repair
  • C0 and A 17 vacuum work
  • F17 kicker replacement
  • Dehumidifier work
  • Vacuum pump maintenance
  • Cryo maintenance
  • Bad stand replacement?

SwitchYard

They are moving SWICs and two quadrupole magnets

NuMI

MI-65 is considered a construction zone. You need steel-toed shoes, long pants, and a hard hat to enter. There is new crane electronics.

MiniBooNE

  • Reverse Horn Polarity
    • Take wipes.
    • Reverse polarity of horn power supply.
    • Test horn.
  • Check Corrosion Tabs at 25m Absorber
    • Unstack 9 blocks from 25m absorber.
    • Retrieve corrosion tabs.
    • Restack blocks.
    • Seal cover.
  • Dry out horn (09Aug07)
  • Check Leakage Current on Ion Pumps
  • Power Supply Maintenance
  • Toroid Work

MTA Beamline Installation

  • Linac shutdown does not start until next week
    • However we had the chance to access the Linac enclosure 3 days this week and complete some critical tasks
      • Alignment
        • Finished setting up controls
        • Beam line stake out is complete
    • Removal of the “gravel” ramp is complete
    • Varieties of small tours with contractors and task managers to assess jobs and confirmation of drawings were complete
  • MTA beam line installation is scheduled to start its tunnel activities on Tuesday, August 14th
    • Equipment readiness
      • All magnets are staged at IB2, on their stands, ready for installation
      • Ship to Linac parking lot next Tuesday, August 14th
  • Next Week:
    • 150 ton mobile crane will be located at south of Linac parking lot for ~ 2 ½ weeks, starting August 14th
      • This is for the removal of the shielding of the MTA hatch and installation of the beam line.
    • The area will also be used as staging area for the needed additional shielding and temporary storage of the magnets and components that will be installed in the tunnel
      • The components are scheduled to be showing up to the area on Monday, August 13th
    • As a result, about 20 parking spaces will be used

Collider Detector at Fermilab

Alignment has conducted as found surveys. They will be opening up their detector.

D-Zero

Alignment has conducted as found surveys. They are conducting system work.

Alignment

Alignment work has begun

Mechanical

Janitorial work in the tunnels has begun

Cryogenic System

  • Shutdown started Sunday morning with fast warm-ups of C4 and D1 magnet strings to room temperature.
    • D1 ODH event occurred when Tevatron Cryogenic Transfer Line LHe supply valve EVSH didn’t properly seal thus venting helium into FRIG bldg; resolved by closing valve with auxiliary control box.
  • Full Tevatron warm-up to LN2 temperature was started Sunday evening after KRSS power outages.
    • 90% of the helium inventory recovered into CHL LHe Dewars and Tank Farm.
    • Tevatron Cryogenic Transfer Line vacuum problems at F2-F3, also at various Pbar components/transfer lines; all vacuums pumped down. Tunnel magnet cryostat vacuum problems being addressed by MSD.
    • Power outage of MSS on Monday morning uneventful.

Cryo Operations

  • 3 FRIG Mycom compressors (plus DZero) running with helium pre-cooled to LN2 temperature flowing through magnet strings
  • CHL helium inventory management system operating (Tank Farm, Helium Purifier, A0 GHe Recovery)
  • CHL Nitrogen Reliquefier plant off and LN2 supply dependent on vendor deliveries to CHL LN2 Dewar tanks
  • Just-in-time shutdown due to CHL feeder 38 failure during the night.

FRIG Maintenance

  • 7 out of 13 wet expander flywheels have been replaced
  • Wet expander overhauls of cold ends have started
  • CDF BA#1 compressor unit is being replaced (failed last month)
  • Various vacuum, instrumentation, and controls problems being worked on.

CHL Maintenance

  • Helium Coldbox-2 was cooled down successfully for readiness testing
  • Coldbox-1
    • Being warmed up to room temperature, as is Coldbox-2
    • Charcoal Adsorbers being set up for high temperature regeneration
  • Nitrogen Reliquefier plant maintenance in progress:
    • Marley Cooling Tower cleanup/overhaul,
    • Rotoflow compressor after-cooler heat exchanger tube bundle being replaced
  • Various facility maintenance to commence next week once CHL operations shiftwork ceases this Friday.

Controls

This Week

  • Recovery from Power Outages
  • Networks, CAMAC Links, Front ends, etc.
  • General Maintenance
  • CAMAC Front End SLD Updates Started
  • Tevatron Copper BP/ABT link decommissioned
  • Tevatron Copper MDAT link decommissioned
  • Main Injector C200 Updates completed

Next Week

  • Saturday 8/11
    • AD Server Updates; loss of access to individual servers as updated. Web access interruptions.
  • Monday 8/13
    • Database Table Update; no DABBEL access ~2hr.
  • Tuesday 8/14
    • MAC Room & AD Computer Room UPS Maintenance
    • Network Switch IOS Updates
  • Wednesday 8/15
    • Network Switch IOS Updates
    • JAVA Engine Restarts/Updates
  • Thursday 8/16
    • Network Switch IOS Updates

Safety

  • Several safety concerns and suggestions reported this week.
  • Numerous requests for PPE, LOTO locks, tags, and job lock boxes
    • Baseball style bump caps are available from the ESH Department for head protection in non-construction areas.
  • MI-65 has been posted as a construction area
    • Hard hats, long pants, and short sleeve shirt are necessary for entering the building or accessing the Pre-Target and Target Hall enclosures.
  • ALARA job estimates are currently coming in according to plans.
  • Utilize Time, Distance, and Shielding to minimize exposure.
  • Task Managers and Crew Leaders are reminded to turn in weekly dosimeter reports to the ESH Department on Monday for the this week.
  • A successful shutdown is getting our work done safely.

FESS

No Report


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/10/2007    email Fermilab